# Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method

### v2

@@ Abstract
In the Venture Capital(VC) industry, predicting the success of startups is challenging due to limited financial data and the need for subjective revenue forecasts. Previous methods based on time series analysis or deep learning often fall short as they fail to incorporate crucial inter-company relationships such as competition and collaboration. Regarding the issues, we propose a novel approach using GrahphRAG augmented time series model. With GraphRAG, time series predictive methods are enhanced by integrating these vital relationships into the analysis framework, allowing for a more dynamic understanding of the startup ecosystem in venture capital. Our experimental results demonstrate that our model significantly outperforms previous models in startup success predictions. To the best of our knowledge, our work is the first application work of GraphRAG.
Introduction
@@
Startups typically represent newly established business models associated with disruptive innovation and high scalability. They are often viewed as powerful engines for economic and social development. With the ongoing economic and social development and the continuous emergence of new technologies, the number of startups has surged, making the prediction of new technologies or business models highly uncertain. Venture capital (VC) evaluation involves assessing investment opportunities in startups and early-stage companies, requiring effective decision-making and data analysis. Valuation is crucial for VCs when deciding to invest in a startup, yet nearly 50% of early-stage VCs make intuitive investment decisions using qualitative methods like gut feelings (Gompers et al. 2020). To avoid relying solely on human expertise and intuition, investors often adopt data-driven approaches to predict the success probability of startups.
@@
Previous work has primarily focused on simple deep learning and machine learning methods, which often rely on limited financial data or single time series analyses. Specifically, most models focus on using historical financial indicators, market trends, or other single-dimensional data to predict the future performance of startups. While these models can capture the growth potential of startups in some cases, they often overlook critical factors within the broader ecosystem in which startups operate, especially the competitive and cooperative relationships between companies. Startups typically operate in environments with limited resources, and their interactions with one another play a significant role in their success or failure. Traditional methods have failed to fully leverage this relational data, leading to poor performance in predicting startup success, particularly in cases with sparse data and low signal strength. These models struggle to accurately identify early success signals under such conditions.
@@
To address these challenges, we propose a novel approach that combines GraphRAG technology with multivariate time series analysis to improve the prediction of startup success. The core of GraphRAG technology lies in its ability to effectively integrate information from various sources, particularly through knowledge graphs and retrieval-augmented generation models, to deeply explore the complex network of relationships among startups. Knowledge graphs can showcase collaborations, competition, supply chain relationships between companies, and other critical information essential for understanding a company’s positioning within its industry ecosystem. By combining this relational information with multivariate time series data, we can build more comprehensive and sophisticated prediction models, significantly improving prediction accuracy and performance on companies with sparse data. Specifically, traditional time series analysis methods often struggle to provide sufficient information support for startups with sparse data and limited samples, leading to instability and low accuracy in prediction results. However, GraphRAG technology enhances model performance by integrating multi-source data, particularly incorporating relationship information among companies into the prediction models, allowing for supplementation and enhancement even in sparse data conditions.
@@
Startups have various potential development goals, with the most desirable being successful acquisition by a large company or going public. While many studies and VCs focus on predicting whether a company will successfully go public as a binary classification task, we attempt to predict what will happen to the company next. This approach aims to provide VCs with opportunities not only to consider high-risk, high-reward companies but also to build lower-risk investment portfolios and determine when to exit.
@@
This paper utilizes two publicly available Chinese news text datasets, selecting data and unstructured news text from companies listed on the A-share market in China from 2013 to 2024. We extracted news reports related to A-share listed companies from these datasets, combined with these companies’ financial data, market performance, and other relevant time series data to construct a comprehensive dataset. Various quantitative and qualitative factors influence the success of startups. In constructing the dataset, we not only collected news reports related to A-share listed companies for the specified years but also conducted an in-depth analysis of the various quantitative and qualitative factors mentioned in these reports. These factors include but are not limited to changes in company executives, industry dynamics, market competition, intellectual property (such as patents, trademarks, etc.). These news reports were transformed into quantifiable indicators using large models for Named Entity Recognition (NER) and structured output, allowing for effective integration with financial data, market performance, and other structured data. By integrating these quantitative and qualitative factors, we strive to build a predictive model that comprehensively reflects the potential of startups.
Our main contributions are as follows:
@@
    We propose a GraphRAG-enhanced multivariate seq2seq time series analysis method to predict startup success in the context of venture capital (VC). As far as we know, our work is the first application work of GraphRAG.
@@
    We review previous related work and identify the lack of crucial relationship information, such as collaboration and competition, in existing methods. We then propose using GraphRAG to extract these relationships from unstructured text data and integrate them into the multivariate time series analysis framework.
@@
    We define startup success prediction as a multivariate seq2seq time series analysis task and conduct extensive experiments on multiple datasets, comparing our method with widely adopted baselines and demonstrating its effectiveness using various evaluation metrics.
The remainder of this paper is organized as follows: Related Work section reviews the related literature, focusing on startup success prediction, venture capital analysis, and the application of knowledge graphs and multivariate time series models. We analyze the limitations of existing prediction models and discuss recent research progress in these fields and their potential in venture capital. Proposed method section details our methodology. Experiment introduces the data collection and processing process, and discusses the experimental setup and results of model comparisons. Finally, the last sections summarize our contributions and propose limitations and future research directions.
Related Work
@@
Figure 1: Our proposed method framework overview: The upper part of the framework represents the process of extracting a knowledge graph from structured text using GraphRAG, followed by transforming it into a mask matrix using the Leiden algorithm. The lower part represents our multivariate sequence-to-sequence LSTM model (Mou et al. 2022).
Deep Learning/Machine Learning Methods
@@
Predicting the success of startups has garnered increasing attention from scholars, especially in the venture capital field. Existing research primarily focuses on developing methods to represent company features and relationships and on building predictive models using machine learning and deep learning techniques. For example, (Cao et al. 2024) constructed CompanyKG to represent and learn various features and relationships between companies, aiming to capture the impact of these features on a company’s success. (Żbikowski and Antosiuk 2021) used Crunchbase data to create an XGBoost-based model for predicting the likelihood of a company’s success. (Kim, Kim, and Sohn 2020) used a doc2vec deep learning method to extract feature vectors from startup profiles and patent abstracts, applying them to filter promising startups in certain fields.In addition, (Ross et al. 2021) introduced the CapitalVX framework, which leverages general business information of startups to predict optimal entry and exit timings, providing valuable insights into market dynamics and investment strategies. Although these approaches have made some progress, they still face challenges in handling complex nonlinear relationships and long-term predictions. (Lyu et al. 2021) designed an incremental representation learning mechanism and a sequential learning model based on graph neural networks to predict the success of startups. (Wang, Zheng, and Wu 2020) performed preprocessing and exploratory analysis on Kickstarter data and then introduced deep learning algorithms like multilayer perceptrons for predicting crowdfunding outcomes. To address long-term revenue prediction, (Cao et al. 2022) proposed the Simulation of Information Revenue Extrapolation (SiRE) algorithm to generate fine-grained long-term revenue forecasts.Furthermore, some studies, such as (Leogrande, Costantiello, and Laureti 2021), have explored the use of unstructured data, including news articles and social media content, for predicting startup success, highlighting the potential of diverse data sources in this domain.
@@
Retrieval-Augmented Generation
In the finance domain, Retrieval-Augmented Generation (Lewis et al. 2021) has become a significant method, particularly for answering questions based on large datasets. RAG methods extract information from retrievable text regions, providing a solid foundation for generation tasks. For instance, (Wang and Ihlamur 2024) developed a startup success prediction framework (SSFF) that incorporates RAG as an external knowledge module to gain a comprehensive understanding of market landscapes. However, in more complex scenarios, such as handling multi-dimensional data and intricate relationships, the capabilities of traditional RAG may be limited.
To address this issue, GraphRAG has been introduced as a structured, hierarchical retrieval-augmented generation method. It not only relies on semantic search of plain text fragments but also enhances the model’s reasoning ability by constructing and utilizing knowledge graphs (Edge et al. 2024). The GraphRAG process involves extracting information from all kinds of unstructured text data (e.g., PDFs, CSVs), building knowledge graphs, generating hierarchical summaries of data, and using the Leiden algorithm to detect industry communities within the knowledge graph (Traag, Waltman, and van Eck 2019). This process aids in intuitively understanding the collaborative and competitive relationships between companies when dealing with complex, multi-dimensional data.
@@
Proposed Method
Our approach consists of two stages: First, we input a large amount of unstructured news text into GraphRAG, where during the clustering and retrieval process, global statistics are dynamically updated to accurately cluster relationships between entities. This results in a directed knowledge graph of companies, with edges representing both direct and indirect relationships between companies. In the second stage, inspired by (Ibrahim et al. 2022) and (Barigozzi and Brownlees 2019), we transform the knowledge graph into a mask matrix using the Leiden algorithm (Traag, Waltman, and van Eck 2019). This matrix is then used as a regularizer in the multivariate time series model, combined with relevant covariates. This method effectively enhances the generalization ability of models on scarce company data and significantly improves prediction performance, as illustrated in Figure 1. The specific algorithm details are provided in Algorithms 1 and 2.
Convolutional LSTM Modeling
Recall the Long Short Term Memory RNN(LSTM) originally proposed by (Hochreiter and Schmidhuber 1997), every hidden layers in the LSTM network are fully connected layers. Formally,
$\begin{array}[]{rl}&i_{t}=\sigma(W_{i}*[X_{t},H_{t-1}]+b_{i})\\ &f_{t}=\sigma(W_{f}*[X_{t},H_{t-1}]+b_{f})\\ &o_{t}=\sigma(W_{o}*[X_{t},H_{t-1}]+b_{o})\\ &C_{t}=f_{t}\circ C_{t-1}+i_{t}\circ\tanh(W_{c}*[X_{t},H_{t-1}]+b_{c})\\ &H_{t}=o_{t}\circ\tanh(C_{t})\end{array}$
where $X_{t}$, $H_{t}$, $C_{t}$ represent the input data, hidden state, and cell state at time step $t$; $i_{t}$, $f_{t}$, $o_{t}$ represent the output, forget, and output gates at time step $t$; $W_{i}$, $W_{f}$, $W_{o}$ represent the convolutional kernel of the input, forget, and output gate; $b_{i}$, $b_{f}$, $b_{o}$ represent the bias of the input, forget, and output gate which is an encoding of everything the cell has observed until time t. The convolution operator is denoted by * and $\circ$ denotes the point-wise product operation. Parameter sharing is a critical aspect for enhancing the generalization capabilities of deep learning models is achieved through the convolutional operations between the kernels and the input data.
The objective of the LSTM is to estimate the conditional probability $p(y_{1},\ldots,y_{T^{\prime}}|x_{1},\ldots,x_{T})$, where $x_{1},\ldots,x_{T}$ represents the input sequence and $y_{1},\ldots,y_{T^{\prime}}$ is its corresponding output sequence, with $T^{\prime}$ possibly differing from $T$. The LSTM computes this conditional probability by first obtaining the fixed-dimensional representation $v$ of the input sequence $x_{1},\ldots,x_{T}$ given by the last hidden state of the LSTM. Subsequently, it computes the probability of $y_{1},\ldots,y_{T^{\prime}}$ using a standard LSTM language model formulation, where the initial hidden state is initialized with the representation $v$ of $x_{1},\ldots,x_{T}$:
$p(y_{1},\ldots,y_{T^{\prime}}|x_{1},\ldots,x_{T})=\prod_{t=1}^{T^{\prime}}p(y_{t}|v,y_{1},\ldots,y_{t-1})$
where each $p(y_{t}\mid v,y_{1},\ldots,y_{t-1})$ distribution is represented with a softmax over all possible feature combinations.
Sequence-to-Sequence Representation
@@
The output sequence $Y_{n}$ represents the company’s post-IPO performance over several quarters, specifically in terms of the price-to-book ratio (P/B ratio). The P/B ratio, defined as the ratio of the company’s stock price (Price) to its book value (Book Value), serves as a crucial metric in venture capital. Unlike the price-to-earnings ratio (P/E ratio), which can be volatile in the short term due to fluctuations in earnings, the P/B ratio offers a more stable measure of a company’s intrinsic value and long-term sustainability (Nissim and Penman 2003). By predicting the P/B ratio over multiple quarters, our model provides a detailed projection of a company’s ongoing financial health and shareholder profitability.
@@
Figure 2: The impact of different mask matrix regularization strengths $\lambda$ on the model’s prediction performance (R-squared).
@@
Figure 3: The knowledge graph generated using GraphRAG and the Leiden algorithm, illustrated using OpenORD (Martin et al. 2011), different colors represent different industries (community clusters).
Leiden Algorithm
Let $G(V,E,w)$ be an undirected graph where $V$ represents the vertex set, $E$ represents the edge set, and $w_{ij}=w_{ji}$ represents the weight of the edge between vertices $i$ and $j$. For unweighted graphs, each edge is assumed to have a unit weight, i.e., $w_{ij}=1$. The set of neighbors for a vertex $i$ is defined as $J_{i}=\{j\ |\ (i,j)\in E\}$. The weighted degree of a vertex $i$ is given by $K_{i}=\sum_{j\in J_{i}}w_{ij}$. The total number of vertices is denoted by $N=|V|$, the total number of edges by $M=|E|$, and the sum of all edge weights in the graph is $m=\sum_{i,j\in V}w_{ij}/2$.
Disjoint community detection focuses on determining a mapping of community membership, represented as $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is associated with a community identifier $c$ from the set of community identifiers $\Gamma$. The vertices belonging to a community $c\in\Gamma$ are denoted by $V_{c}$, and the community to which a vertex $i$ belongs is represented by $C_{i}$. Additionally, the neighbors of vertex $i$ that are part of community $c$ are denoted as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, with the total weight of these edges given by $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$. The total weight of edges within community $c$ is represented by $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the overall edge weight associated with community $c$ is denoted by $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec 2021).
Modularity is a measure used to assess the quality of communities identified by heuristic-based community detection algorithms. It is determined by calculating the difference between the actual fraction of edges within communities and the expected fraction of such edges if they were randomly distributed. The resulting value typically falls within the range of $[-0.5,1]$, where higher values signify superior results (Brandes et al. 2007). The optimization of this metric theoretically leads to the optimal grouping (Newman 2004). The modularity $Q$ of identified communities is determined using modularity equation, where delta represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise). The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, can be calculated as follows:
$Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})$
$\Delta Q_{i:d\rightarrow c}=\frac{1}{m}(K_{i\rightarrow c}-K_{i\rightarrow d})-\frac{K_{i}}{2m^{2}}(K_{i}+\Sigma_{c}-\Sigma_{d})$

### v4

@@ Abstract
In the Venture Capital (VC) industry, predicting the success of startups is challenging due to limited financial data and the need for subjective revenue forecasts. Previous methods based on time series analysis often fall short as they fail to incorporate crucial inter-company relationships such as competition and collaboration. To fill the gap, this paper aims to introduce a novel approach using GraphRAG augmented time series model. With GraphRAG, time series predictive methods are enhanced by integrating these vital relationships into the analysis framework, allowing for a more dynamic understanding of the startup ecosystem in venture capital. Our experimental results demonstrate that our model significantly outperforms previous models in startup success predictions.
Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method
Zitian Gao1, Yihao Xiao2 1School of Computer Science, The University of Sydney, Australia 2School of Finance, Shanghai University of Finance and Economics, China zgao3186@uni.sydney.edu.au, 2023121716@stu.sufe.edu.cn
1 Introduction
@@
Startups typically represent newly established business models associated with disruptive innovation and high scalability. They are often viewed as powerful engines for economic and social development Cao et al. (2023). With the ongoing economic and social development and the continuous emergence of new technologies, the number of startups has surged, making the prediction of new technologies or business models highly uncertain. Venture capital (VC) evaluation involves assessing investment opportunities in startups and early-stage companies, requiring effective decision-making and data analysis. Valuation is crucial for VCs when deciding to invest in a startup, yet nearly 50% of early-stage VCs make intuitive investment decisions using qualitative methods like gut feelings Gompers et al. (2020). To avoid relying solely on human expertise and intuition, investors often adopt data-driven approaches to predict the success probability of startups Corea et al. (2021); Hunter et al. (2018).
@@
Previous work has primarily focused on simple deep learning and machine learning methods, which often rely on limited financial data or single time series analyses Wang and Ihlamur (2024); Lewis et al. (2021). Specifically, most models focus on using historical financial indicators, market trends, or other single-dimensional data to predict the future performance of startups. While these models can capture the growth potential of startups in some cases, they often overlook critical factors within the broader ecosystem in which startups operate, especially the competitive and cooperative relationships between companies. Startups typically operate in environments with limited resources, and their interactions with one another play a significant role in their success or failure. Traditional methods have failed to fully leverage this relational data, leading to poor performance in predicting startup success, particularly in cases with sparse data and low signal strength. These models struggle to accurately identify early success signals under such conditions.
@@
To address these challenges, we propose a novel approach that combines GraphRAG technology with multivariate time series analysis to improve the prediction of startup success. The core of GraphRAG technology lies in its ability to effectively integrate information from various sources, particularly through knowledge graphs and retrieval-augmented generation models Edge et al. (2024), to deeply explore the complex network of relationships among startups. Knowledge graphs can showcase collaborations, competition, supply chain relationships between companies, and other critical information essential for understanding a company’s positioning within its industry ecosystem. By combining this relational information with multivariate time series data, we can build more comprehensive and sophisticated prediction models, significantly improving prediction accuracy and performance on companies with sparse data. Specifically, traditional time series analysis methods often struggle to provide sufficient information support for startups with sparse data and limited samples, leading to instability and low accuracy in prediction results. However, GraphRAG technology enhances model performance by integrating multi-source data, particularly incorporating relationship information among companies into the prediction models, allowing for supplementation and enhancement even in sparse data conditions.
@@
Startups have various potential development goals, with the most desirable being successful acquisition by a large company or going public. While many studies and VCs focus on predicting whether a company will successfully go public as a binary classification task Cao et al. (2022); Wang and Ihlamur (2024), we attempt to predict what will happen to the company next. This approach aims to provide VCs with opportunities not only to consider high-risk, high-reward companies but also to build lower-risk investment portfolios and determine when to exit Bienz and Walz (2010); Schwienbacher (2008).
@@
This paper utilizes two publicly available Chinese news text datasets He et al. (2023); Lu et al. (2023), selecting data and unstructured news text from companies listed on the A-share market in China from 2013 to 2024. We extracted news reports related to A-share listed companies from these datasets, combined with these companies’ financial data, market performance, and other relevant time series data to construct a comprehensive dataset. Various quantitative and qualitative factors influence the success of startups. In constructing the dataset, we not only collected news reports related to A-share listed companies for the specified years but also conducted an in-depth analysis of the various quantitative and qualitative factors mentioned in these reports. These factors include but are not limited to changes in company executives, industry dynamics, market competition, intellectual property (such as patents, trademarks, etc.). These news reports were transformed into quantifiable indicators using large models for Named Entity Recognition (NER) and structured output, allowing for effective integration with financial data, market performance, and other structured data. By integrating these quantitative and qualitative factors, we strive to build a predictive model that comprehensively reflects the potential of startups.
In contrast to another common direction in financial time series prediction—stock trend forecasting—fine-grained, structured time series data used for predicting stock trends is readily available Sawhney et al. (2021); Wang et al. (2021, 2022). By comparison, predicting the success of startups in the venture capital (VC) field requires forecasting the success of unlisted startups with minimal available data, often relying on hard-to-access unstructured text such as team backgrounds, patents, and product reviews.
Stock trend forecasting typically utilizes fine-grained data at the second or minute level, whereas startup success prediction in the VC domain often relies on coarse-grained data at the quarterly or annual level. Additionally, relationships such as collaboration, competition, and investment activity play a more significant role in the VC industry compared to stock forecasting in quantitative finance. These distinctions pose unique challenges for predicting startup success in the VC field Baierl et al. (2016); Pahnke et al. (2014).
To address these issues, we proposed our GraphRAG augmented multivariate time series method. Our main contributions are as follows:
@@
    We propose a GraphRAG-enhanced multivariate Seq2Seq time series analysis method to predict startup success in the context of venture capital (VC).
@@
    We review previous related work and identify the lack of crucial relationship information, such as collaboration and competition, in existing methods. We propose using GraphRAG to extract these relationships from unstructured text data and integrate them into the multivariate time series analysis framework.
@@
    We introduce our GraphRAG-based data pipeline, which extracts structured information from extensive news text to address the venture capital industry’s data scarcity for multivariate time series analysis. We then present a novel method to integrate key knowledge graph information, significantly improving startup success prediction performance.
2 Related Work
@@
Figure 1: Our proposed method framework overview: The upper part of the framework represents the process of extracting a knowledge graph from structured text using GraphRAG, followed by transforming it into a mask matrix using the Leiden algorithm. The lower part represents our multivariate sequence-to-sequence LSTM model Mou et al. (2022).
2.1 Startup Success Prediction
@@
Predicting the success of startups has garnered increasing attention from scholars, especially in the venture capital field. Existing research primarily focuses on developing methods to represent company features and relationships and on building predictive models using machine learning and deep learning techniques. For example, Cao et al. (2024) constructed CompanyKG to represent and learn various features and relationships between companies, aiming to capture the impact of these features on a company’s success. Żbikowski and Antosiuk (2021) used Crunchbase data to create an XGBoost-based model for predicting the likelihood of a company’s success. Kim et al. (2020) used a doc2vec deep learning method to extract feature vectors from startup profiles and patent abstracts, applying them to filter promising startups in certain fields.In addition, Ross et al. (2021) introduced the CapitalVX framework, which leverages general business information of startups to predict optimal entry and exit timings, providing valuable insights into market dynamics and investment strategies. Although these approaches have made some progress, they still face challenges in handling complex nonlinear relationships and long-term predictions. Lyu et al. (2021) designed an incremental representation learning mechanism and a sequential learning model based on graph neural networks to predict the success of startups. Wang et al. (2020) performed preprocessing and exploratory analysis on Kickstarter data and then introduced deep learning algorithms like multilayer perceptrons for predicting crowdfunding outcomes. To address long-term revenue prediction, Cao et al. (2022) proposed the Simulation of Information Revenue Extrapolation (SiRE) algorithm to generate fine-grained long-term revenue forecasts.Furthermore, some studies, such as Leogrande et al. (2021), have explored the use of unstructured data, including news articles and social media content, for predicting startup success, highlighting the potential of diverse data sources in this domain.
@@
2.2 Retrieval-Augmented Generation In Finance
In the finance domain, Retrieval-Augmented Generation Lewis et al. (2021) has become a significant method, particularly for answering questions based on large datasets. RAG methods extract information from retrievable text regions, providing a solid foundation for generation tasks. For instance, Wang and Ihlamur (2024) developed a startup success prediction framework (SSFF) that incorporates RAG as an external knowledge module to gain a comprehensive understanding of market landscapes. However, in more complex scenarios, such as handling multi-dimensional data and intricate relationships, the capabilities of traditional RAG may be limited.
To address this issue, GraphRAG has been introduced as a structured, hierarchical retrieval-augmented generation method. It not only relies on semantic search of plain text fragments but also enhances the model’s reasoning ability by constructing and utilizing knowledge graphs Edge et al. (2024). The GraphRAG process involves extracting information from all kinds of unstructured text data (e.g., PDFs, CSVs), building knowledge graphs, generating hierarchical summaries of data, and using the Leiden algorithm to detect industry communities within the knowledge graph Traag et al. (2019). This process aids in intuitively understanding the collaborative and competitive relationships between companies when dealing with complex, multi-dimensional data.
@@ 2.2 Retrieval-Augmented Generation In Finance
3 Methodology
Our approach consists of two stages: First, we input a large amount of unstructured news text into GraphRAG, where during the clustering and retrieval process, global statistics are dynamically updated to accurately cluster relationships between entities. This results in a directed knowledge graph of companies, with edges representing both direct and indirect relationships between companies. In the second stage, inspired by Ibrahim et al. (2022) and Barigozzi and Brownlees (2019), we transform the knowledge graph into a mask matrix using the Leiden algorithm Traag et al. (2019). This matrix is then used as a regularizer in the multivariate time series model, combined with relevant covariates. This method effectively enhances the generalization ability of models on scarce company data and significantly improves prediction performance, as illustrated in Figure 1. The specific algorithm details are provided in Algorithms 1 and 2.
3.1 Sequence-to-Sequence Representation
@@
The output sequence $Y_{n}$ represents the company’s post-IPO performance over several quarters, specifically in terms of the price-to-book ratio (P/B ratio). The P/B ratio, defined as the ratio of the company’s stock price (Price) to its book value (Book Value), serves as a crucial metric in venture capital. Unlike the price-to-earnings ratio (P/E ratio), which can be volatile in the short term due to fluctuations in earnings, the P/B ratio offers a more stable measure of a company’s intrinsic value and long-term sustainability Nissim and Penman (2003). By predicting the P/B ratio over multiple quarters, our model provides a detailed projection of a company’s ongoing financial health and shareholder profitability.
@@
Figure 2: The knowledge graph generated using GraphRAG and the Leiden algorithm, illustrated using OpenORD Martin et al. (2011), different colors represent different industries (community clusters).
3.2 Graph Clustering
Let $G(V,E)$ be an undirected graph where $n:=|V|$ represents the number of vertices, and $m:=|E|$ represents the number of edges. We define $C=\{C_{1},\dots,C_{k}\}$ as a partition of the vertex set $V$. The set $C$ is referred to as a clustering of $G$, with each $C_{i}$ being non-empty. A clustering $C$ is termed trivial if $k=1$ or $k=n$. The collection of all possible clusterings of $G$ is denoted by $\mathcal{A}(G)$. In what follows, a cluster $C_{i}$ is frequently associated with the induced subgraph of $G$, denoted as $G[C_{i}]:=(C_{i},E(C_{i}))$, where
$E(C_{i}):=\{\{v,w\}\in E:v,w\in C_{i}\}.$
Thus, we define $E(C):=\bigcup_{i=1}^{k}E(C_{i})$ as the collection of intracluster edges, while the complement set $E\setminus E(C)$ represents the intercluster edges. The total number of intracluster edges is denoted by $m(C)$ with the total number of intercluster edges by $\overline{m}(C)$. Additionally, the set of edges that connect a node in $C_{i}$ to a node in $C_{j}$ is denoted by $E(C_{i},C_{j})$.
In complex networks, nodes tend to group together, forming relatively dense clusters commonly referred to as communities. Modularity is a measure used to assess the quality of communities identified by heuristic-based community detection algorithms. The formulation of modularity from Brandes et al. (2008) is as follows:
${q}({C}):=$ $\sum_{C\in{C}}\left[{{|E({C})|}\over{m}}-\left({{|E({C})|+\sum_{C^{\prime}\in{C}}|E(C,C^{\prime})|}\over{2m}}\right)^{2}\right].$
The Leiden algorithm originally introduced by Traag et al. (2019) is a multi-level approach to community detection aiming to maximizing the modularity gain. This algorithm involves three stages. The first stage is known as the local-moving phase, involves each vertex $i$ optimizing its community assignment by greedily selecting to join the community of one of its neighbors denoted as $j\in J_{i}$. The next stage is the refinement phase, where vertices within each community further update their community assignments, starting from singleton partitions. Unlike the local-moving phase, these updates are not purely greedy; instead, vertices may move to any community within their scope where the modularity gain is positive, with the likelihood of moving to a neighboring community proportional to the modularity gain of that move. Finally, the aggregation phase combines all vertices within each refined partition into super-vertices, with an initial community assignment inherited from the local-moving phase Traag et al. (2019).